# Quasicrystals as an intermediate form of matter between crystalline and amorphous solids

### v1

@@ II.3 Quasicrystal structure and dynamics
These two quantities, $g\left(r\right)$ and $S\left(k\right)$, play pivotal roles in characterizing the structure of phases from two complementary perspectives.
Furthermore, the mean-squared displacement (MSD) serves as a standard method to probe the dynamics of materials. It typically quantifies the average displacement of a particle within the system as a function of time, expressed as,
@@
However, in 1D and 2D systems, long-range thermal fluctuations in positional ordering, as described by the Mermin-Wagner fluctuations [81], can pose challenges for accurately measuring real dynamics. These fluctuations manifest as appreciable effects, as the system size imposes a cutoff on the maximum wavelength [82, 83, 84].
@@
Mitigating the finite-size effects when studying dynamics in the 2D dodecagonal quasicrystal system is especially important when using periodic boundary conditions. To address this problem, we also employ a modified method by calculating the “cage-relative MSD” (CR-MSD) in place of the standard MSD:
$\left\langle r^{2}\left(t\right)\right\rangle_{\rm{CR}}=\frac{1}{N}\left\langle\sum_{i=1}^{N}\left[\left(\mathbf{r}_{i}\left(t\right)-\mathbf{r}_{i}\left(0\right)\right)-\frac{1}{N_{j}}\sum_{j=1}^{N_{j}}\left(\mathbf{r}_{j}\left(t\right)-\mathbf{r}_{j}\left(0\right)\right)\right]^{2}\right\rangle.$ (7)
In this context, the second term in Eq. (7) represents the center of mass of the cage, which is derived from the positions of the $N_{j}$ nearest neighbors of the central particle $i$. This approach helps mitigating finite-size effects and allows for a more accurate assessment of dynamics in the 2D dodecagonal quasicrystal system.
In most cases, particles exhibit ballistic motion at very short times and eventually transition to random diffusive motion at longer timescales. As such, Mean square displacement (MSD), or the cage-relative MSD (CR-MSD in the context of a 2D system), can be employed to assess the diffusivity $D$ of a $d$-dimensional system, utilizing the Einstein relation:
@@
The determination of this quantity provides an important point of contact between experiments on quasi-crystalline materials and our simulations.
@@
where $\Delta\mathbf{r}_{i}\left(t\right)=\mathbf{r}_{i}\left(t+\tau\right)-\mathbf{r}_{i}\left(\tau\right)$, with $\tau$ representing a time interval. The wavenumber $k$ is chosen as the first peak of the static structure factor $S(k)$, which in the present system this corresponds to $k^{*}$. The structural relaxation time $\tau_{\alpha}$ is then defined as the time when $F_{\text{s}}\left(k^{*},\tau{\alpha}\right)=e^{-1}$.
@@
This approach allows us to account for nearest neighbor contributions and effectively characterize the dynamics in the presence of structural constraints.
@@
However, in line with the approach employed for the cage-relative MSD and SISF, we precisely calculate the cage-relative non-Gaussian parameter, which is defined as:

### v2

@@ II.2 Phase diagram
We generate the intensity maps of the static structure factor using Eq. (4), which provide a comprehensive representation of the 2D structures included in the phase diagram. These intensity maps resemble the non-energy-resolved diffraction patterns typically obtained in diffraction experiments to determine crystal structure and symmetry. The diffraction pattern of the disordered phase (Fig. 4(a)) consists of diffuse rings of light, resulting from the lack of regular arrangement in ordered lattices. In contrast, the diffraction pattern of the hexatic phase exhibits both weak polyhedra and intense reflections with 6-fold and 12-fold symmetries (Fig. 4(b)), clearly indicating the hexatic phase intermediate nature between the liquid and solid phases. The diffraction pattern of the quasicrystal (Fig. 4(c)) is more intricate, displaying a 12-fold symmetry and well-defined reflections arranged in a quasi-periodic 2D pattern. These diffraction patterns serve to reinforce the credibility of our atomistic model. Note that for each of these diffraction patterns, we conducted parallel simulations on 50 independent configurations to enhance statistical reliability. Further quantification of the average quasicrystal structure will be the focus of our analysis in Appendix A.
We note that there have been many previous simulations of two dimensional crystalline materials in which corresponding transitions from a low temperature hexagonal crystalline state to an orientationally correlated hexatic state occurs upon heating, followed by an additional transition between the hexatic and liquid states at an even higher characteristic temperature. Although the well-known theories of these transitions developed by Kosterlitz and Thouless [56] and Halperin and Nelson [57, 58] indicate that these transitions are continuous, many experimental and computational studies have established that these transitions can be either of second or first order in nature [59], depending on the particular form of the interparticle potential, and the presence of appreciable finite size effects can often make the assignment of transition order computationally challenging and sometimes controversial. Resolving the difference between weakly first order transitions and second order transitions can be inherently difficult [60] and we avoid this technical discussion in the present work in relation to these thermodynamic transitions in quasicrystalline materials.
There have also been instructive studies of the collective dynamics of two dimensional systems exhibiting crystal melting and formation of an intermediate hexatic phase [61, 62, 63]. In the present exploratory study of a model quasicrystalline material, we simply note that the thermodynamic transitions that we observe in our quasicrystal material simulations are strikingly reminiscent to the hexagonal to hexatic and hexatic to liquid transitions established to exist in many two-dimensional crystal forming materials. Moreover, these transitions in our quasicrystalline forming material are relatively sharp in terms of structures in both real and reciprocal space, as well as in energy and heat capacity, suggestive of a weakly first order phase transitions. Given the limited size of our simulated materials, however, we do not insist on this designation of transition order and simply note that apparently well-defined thermodynamic transitions occur near certain characteristic temperatures at which the degree of material order and dynamics clearly occur. We are not aware of a previous work exploring these transitions in quasicrystalline materials.
Previous work on crystal melting have also found a prevalence for string-like collective motion [61, 62], a phenomenon that appears superficially very similar to the observations that we report below in our simulated model quasicrystal forming material. Despite the superficial similarities in the thermodynamic and dynamic properties of two dimensional crystal and quasicrystal forming materials, we stop short of equating their properties. Nonetheless, the evidence of evident long-range positional correlations in our quasicrystalline material point to a greater “similarity” to crystalline materials than glass-forming materials. However, given the somewhat contradictory indications from simulation and measurements of a “similarity” of quasicrystalline materials to both crystalline and glass-forming materials, we take a fresh look at which class of materials more closely resembles quasicrystalline materials from a dynamical perspective.
II.3 Quasicrystal dynamics
The mean-squared displacement (MSD) serves as a standard method to probe the dynamics of materials. The MSD normally quantifies the average displacement of a particle within the system as a function of time, expressed as,
@@ II.3 Quasicrystal dynamics
However, in 1D and 2D systems, long-range thermal fluctuations in positional ordering occurs, as described by the Mermin-Wagner fluctuations [64], can pose challenges for accurately measuring the “real” dynamics of two-dimensional materials. These fluctuations manifest as appreciable effects, as the simulation system size imposes a cutoff on the maximum wavelength of these fluctuations[65, 66, 67]. Mitigating the finite-size effects when studying dynamics in the 2D dodecagonal quasicrystal system is especially important when using periodic boundary conditions. To address this problem, we also employ a modified method by calculating the “cage-relative MSD” (CR-MSD) in place of the standard MSD:
$\left\langle r^{2}\left(t\right)\right\rangle_{\rm{CR}}=\frac{1}{N}\left\langle\sum_{i=1}^{N}\left[\left(\mathbf{r}_{i}\left(t\right)-\mathbf{r}_{i}\left(0\right)\right)-\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}\left(\mathbf{r}_{j}\left(t\right)-\mathbf{r}_{j}\left(0\right)\right)\right]^{2}\right\rangle.$ (6)
In this context, the second term in Eq. (6) represents the center of mass of the cage, which is derived from the positions of the $N_{i}$ nearest-neighbors of the central particle $i$. This approach helps to mitigate finite-size effects and allows for a more accurate assessment of dynamics in the 2D dodecagonal quasicrystal system.
@@
In most cases, particles exhibit ballistic motion at very short times and eventually transition to random diffusive motion at longer timescales. As such, MSD, or the cage-relative MSD (CR-MSD in the context of a 2D system), can be employed to assess the diffusivity $D$ of a $d$-dimensional system, utilizing the Einstein relation:
@@
Here, $d=2$. The determination of this quantity provides an important point of contact between experiments on quasicrystalline materials and our simulations.
@@
where $\Delta\mathbf{r}_{i}\left(t\right)=\mathbf{r}_{i}\left(t+\Delta t\right)-\mathbf{r}_{i}\left(t\right)$, $\Delta t$ represents a time interval. Here, the SISF is defined at the wave vector $k=7.3$, which corresponds to the major peak in $S(k)$ and is sensitive to variations in atomic density, making it an important indicator of structural transitions, particularly in disordered phases. Besides, the temperature dependence of the first peak position in $S(k)$ is weak and the peak position is taken to be a constant. The structural relaxation time $\tau_{\alpha}$ is then defined as the time when $F_{\text{s}}\left(k^{*},\tau_{\alpha}\right)=e^{-1}$.
@@
This approach allows us to account for nearest-neighbor contributions and effectively characterize the dynamics in the presence of structural constraints.
@@
However, in line with the approach employed for the cage-relative MSD and SISF, we precisely calculate the cage-relative non-Gaussian parameter, which is defined as,